2A(g) + B(g) ⇌ C(g)(exothermic)At equilibrium a 2 L container contains 2.0 mol of A, 0.04 mol of B and 1.2 mol of C gases at a constant temperature. If the temperature inside the container is increased and then some of gas A is removed from the container; the molar concentration of B becomes 0.128 M at the new equilibrium and the equilibrium constant becomes 1/10 times of its initial value. What is the number of moles of gas A removed?
1
Expert's answer
2013-03-13T09:25:37-0400
Equilibrium constant is:
K = [C]/[A]^2*[B] = (1.2/2) / (0.04/2) * (2.0/2) = 30
New equilibrium constant is 30/10=3
x 0.5x
2A(g) + B(g) ⇌ C(g)

K = [0.6-0.5x]/[0.02-x]^2*[0.128]=3
x= 0.77 M

real number of moles is 0.77*2=1.54 mol